Randy invested his inheritance in an account that paid 6.5% interest, compounded continuously. After 7 years, he found that he now had $56,449.07. What was the original amount of his inheritance?
Answer by jorel1380(3719)

If its' compounded continuously, then the formula is:
A=Pe^(rt),where P is your starting amount, r is the rate, and t is the time.
------------------
A=Pe^(rt)
A/e^(rt)=P
56449.07/2.7183^(7*.065)=P
56449.07/1.5762="
"""""""=35813.89
--------------
Randy's inheritance was originally worth $35813.89.